OpenStudy (anonymous):

one number is8 more than a second number. The sum of the two number is 18.Find the number..

OpenStudy (anonymous):

Step 1: Let x be that "one number" Step 2: Can you express the second number in terms of x? Step 3: Set up an equation with left side equal to the sum of x and the second number, while the right side being 18 Step 4: Solve it.

OpenStudy (anonymous):

You have 2 number that add up to 18. You can start with 18 and divide by 2 (since there are 2 number). There is an 8 number difference so you need to subtract and add by four to get your two numbers. Your equations are (18/2)-4=x and (18/2)+4=x.

OpenStudy (anonymous):

Excuse me, according to your saying, is (18/2)-4=x OR is (18/2)+4=x?

OpenStudy (anonymous):

@wrstlr3232

OpenStudy (anonymous):

It's both answers. The question is, one number is 8 more than a second number. Wouldn't you need to find both numbers?

OpenStudy (johnweldon1993):

@Armi the question states the second number is 8 MORE than the first...not 8 times

OpenStudy (anonymous):

oops sorry I didn't noticed

OpenStudy (anonymous):

so: x+y=18 x=y+8 y+8+y=18 2y=10 y=5 x=13